# What’s Inside Your Diffusion Model? A Score-Based Riemannian Metric to Explore the Data Manifold

### v2

@@ 4.1 Embedded sphere
We train a diffusion model on the embedded samples and extract the score function $\mathbf{s}(\mathbf{x})$. As shown in Figure 2C, the learned score vectors align with normal vectors of the sphere (quantitative validation in Supplementary Material), replicating the theoretical prediction that score functions are normal to the data manifold [13]. Using these score functions, we compute geodesics and compare with linear interpolation. Figure 2D-F shows that while linear paths cut through the sphere (producing blurred intermediate images), our geodesics follow the manifold’s surface (maintaining high sample quality throughout). Similarly, for extrapolation, our approach follows the sphere’s curvature and preserves image structure, while linear extrapolation quickly departs from the manifold, causing severe distortions.
@@ 4.2 Rotated MNIST
To evaluate whether our score-based metric captures perceptually meaningful transformations, we conducted experiments on a modified version of the MNIST dataset where the transformation structure is well-defined. We created a custom Rotated MNIST dataset by systematically rotating each digit by angles ranging from $0^{\circ}$ to $350^{\circ}$ in $10^{\circ}$ increments, producing 36 rotated versions of each original digit. We then trained a denoising diffusion model on this dataset using a UNet-2D [29] architecture with attention blocks. The model was trained for 100 epochs with a batch size of 256 (see more details in Appendix). Our hypothesis is that this trained model implicitly learns the data manifold, including the rotational structure of the dataset.
For interpolation experiments, we selected test set digit pairs with varying orientations, typically choosing the same digit rendered at different angles. Figure 3A compares our geodesic interpolation method against LERP, SLERP and Noise Diffusion. The results show that our geodesic approach produces trajectories that follow the rotational structure of the data manifold, creating smooth transformations that preserve digit identity while naturally rotating from one orientation to another. While LERP often produces unrealistic blending artifacts where intermediate frames superimpose features from both orientations rather than rotating continuously, our method follows the natural rotation transformation. Quantitatively, we measure the quality of interpolated frames using both PSNR and SSIM [30] metrics (Table 1), with our geodesic approach outperforming other methods in intermediate frame quality.
@@
PSNR $\uparrow$ 14.08 $\pm$ 1.27 13.64 $\pm$ 1.33 13.67 $\pm$ 0.83 14.98 $\pm$ 1.29
SSIM $\uparrow$ 0.578 $\pm$ 0.062 0.572 $\pm$ 0.074 0.568 $\pm$ 0.054 0.650 $\pm$ 0.056
@@ 4.3 Stable Diffusion & MorphBench
To demonstrate our method’s applicability to state-of-the-art diffusion models, we conducted experiments with Stable Diffusion 2.1 [31], a powerful latent diffusion model trained on billions of image-text pairs. This allowed us to scale our approach on a more complex, higher-resolution image generation task than our previous experiments. Stable Diffusion operates in a compressed latent space rather than pixel space, making it an interesting test case for our score-based metric. The diffusion process occurs in a 4×64×64 latent space (representing 512×512 pixel, RGB images), where the data manifold has complex geometry reflecting natural image statistics. We apply our methodology by computing geodesics in this latent space using score estimates from the model’s denoising U-Net, then mapping the results back to image space using the model’s VAE decoder.
For evaluation, we utilize MorphBench [32], a comprehensive benchmark for assessing image morphing capabilities. MorphBench consists of 90 diverse image pairs organized into two categories: (1) metamorphosis between different objects (66 pairs) and (2) animation of the same object with different attributes (24 pairs). This diversity allows us to evaluate our method across varying transformation complexities.
Figure 4 (see Appendix G, for more examples) presents qualitative comparisons between our geodesic method and baseline approaches (LERP, SLERP, and Noise Diffusion). Our method generates smoother, more natural-looking transitions that better preserve image coherence during the morphing process. While LERP and SLERP often produce blurry intermediate frames with unrealistic composites of both source and target images, and Noise Diffusion frequently generates images with artifacts, our geodesic interpolation creates a more natural progression by following meaningful paths on the data manifold.
Table 2 provides a comprehensive quantitative evaluation. Our geodesic method achieves the best performance on LPIPS (0.3582 vs. 0.3613 for LERP), a learned perceptual similarity metric that correlates strongly with human judgments of visual quality [33]. We also outperform all baselines on distribution-level metrics, with the lowest FID (140.61 vs. 148.19 for LERP) [34] and KID scores (0.0863 vs. 0.0935 for LERP) [35]. These results indicate that our interpolated images are both perceptually coherent and maintain high sample quality that better matches the "true" data distribution. While LERP achieves slightly better results on pixel-level metrics (PSNR: 21.28 vs. 20.88; SSIM: 0.6274 vs. 0.6180), it’s well-established that these metrics often fail to capture perceptual quality [33], especially for high-resolution natural images.
@@
avg_SSIM $\uparrow$ 0.627 $\pm$ 0.124 0.575 $\pm$ 0.144 0.404 $\pm$ 0.119 0.618 $\pm$ 0.117
avg_PSNR $\uparrow$ 21.28 $\pm$ 1.81 20.36 $\pm$ 2.25 16.73 $\pm$ 2.03 20.88 $\pm$ 1.78
avg_LPIPS $\downarrow$ 0.361 $\pm$ 0.077 0.396 $\pm$ 0.081 0.480 $\pm$ 0.046 0.358 $\pm$ 0.070
FID $\downarrow$ 148.2 $\pm$ 50.3 171.5 $\pm$ 53.6 271.8 $\pm$ 47.6 140.6 $\pm$ 51.3
KID $\downarrow$ 0.094 $\pm$ 0.052 0.110 $\pm$ 0.057 0.186 $\pm$ 0.064 0.086 $\pm$ 0.055
@@ 5 Discussion
In this work, we introduced a score-based Riemannian metric derived from diffusion models that captures the intrinsic geometry of the data manifold without requiring explicit parameterization. Our approach leverages the Stein score function to define a metric that naturally adapts to the manifold structure, enabling geometric computations that respect the underlying data distribution. Through experiments on synthetic data, MNIST, and complex natural images, we demonstrated that geodesics in our metric space correspond to perceptually meaningful paths between images, outperforming conventional interpolation methods.
@@
An interesting pattern emerged when comparing performance across different datasets. On Rotated MNIST, our method outperformed baselines across pixel-level metrics (PSNR and SSIM). However, on the higher-resolution MorphBench dataset with Stable Diffusion, our method dominated in perceptual metrics (LPIPS) and distribution metrics (FID, KID) but was slightly outperformed by LERP on pixel-level metrics. This difference can be attributed to our geodesic computations for Stable Diffusion occurring in the compressed latent space rather than the pixel space. The VAE decoder introduces additional variability when mapping back to pixels, which may affect pixel-perfect reconstruction while still preserving—and even enhancing—perceptual quality. This observation aligns with the established understanding that pixel-level metrics often fail to capture perceptual similarity in high-resolution natural images, instead favoring blurry but pixel-wise accurate reconstructions over sharper, perceptually preferable images [33].
@@
Limitations. Despite its strengths, our approach introduces additional computational complexity compared to direct methods like LERP or SLERP. Our implementation typically requires several hundred optimization steps to converge, whereas alternative methods can be computed without any iterative optimization procedure.
@@
Another limitation relates to the validation of extrapolation results. Although our experiments demonstrate that the extrapolated images maintain high quality and follow natural extensions of the geodesic paths, there are no established quantitative metrics to evaluate the quality of extrapolations in our settings, making objective comparisons challenging. Nevertheless, we view extrapolation as a data-driven exploration of the learned representation—a tool for discovering the implicit structure captured by diffusion models rather than a task requiring ground truth validation.
@@
Future Work. One exciting direction would be applying our approach to semantic image editing tasks [36, 37]. By computing geodesic paths between edited and original images, our method could enable more natural and realistic transitions when performing attribute manipulations, style transfers, or object insertions.
@@
Additionally, the computational efficiency of geodesic calculation could be improved through techniques like neural surrogate models [38] that directly predict geodesic paths. Such models could enable real-time interactive editing tools that allow artists and designers to explore the learned manifold of a diffusion model while maintaining high sample quality throughout the editing process.
@@
Finally, our score-based metric provides a new lens for analyzing the geometry learned by diffusion models. Future work could explore using this geometric perspective to better understand model behavior, detect biases in learned distributions, or visualize how the data manifold evolves during training. By continuing to develop the connections between diffusion models and Riemannian geometry, we can gain deeper insights into how these powerful generative models represent complex data distributions.
Appendix A Proof of Riemannian Metric Properties
@@
To prove that $g(\boldsymbol{x})=\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T}$ defines a valid Riemannian metric, we need to show that it satisfies the two fundamental properties:
Symmetry: For any point $\boldsymbol{x}$, $g(\boldsymbol{x})$ must be symmetric.
$g(\boldsymbol{x})^{T}$ $=(\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T})^{T}$
$=\mathbf{I}^{T}+\lambda\cdot(\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T})^{T}$
$=\mathbf{I}+\lambda\cdot(\boldsymbol{s}(\boldsymbol{x})^{T})^{T}\boldsymbol{s}(\boldsymbol{x})^{T}$
$=\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T}$
$=g(\boldsymbol{x})$
Positive-definiteness: For any point $\boldsymbol{x}$ and any non-zero vector $\boldsymbol{v}$, we must have $\boldsymbol{v}^{T}g(\boldsymbol{x})\boldsymbol{v}>0$.
$\boldsymbol{v}^{T}g(\boldsymbol{x})\boldsymbol{v}$ $=\boldsymbol{v}^{T}(\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T})\boldsymbol{v}$
$=\boldsymbol{v}^{T}\mathbf{I}\boldsymbol{v}+\lambda\cdot\boldsymbol{v}^{T}\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T}\boldsymbol{v}$
$=\|\boldsymbol{v}\|^{2}+\lambda\cdot(\boldsymbol{s}(\boldsymbol{x})^{T}\boldsymbol{v})^{2}$
Since $\lambda>0$ (by construction) and $\|\boldsymbol{v}\|^{2}>0$ for any non-zero vector $\boldsymbol{v}$, we have $\boldsymbol{v}^{T}g(\boldsymbol{x})\boldsymbol{v}>0$. The second term $\lambda\cdot(\boldsymbol{s}(\boldsymbol{x})^{T}\boldsymbol{v})^{2}$ is always non-negative, further ensuring positive-definiteness.
Thus, the Stein score metric tensor $g(\boldsymbol{x})=\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T}$ satisfies all the requirements of a Riemannian metric on $\mathbb{R}^{N}$.
Appendix B Geodesic Algorithm Details
B.1 Discretization and Numerical Integration

### v3

@@ 4.1 Embedded sphere
We train a diffusion model on the embedded samples and extract the score function $\mathbf{s}(\mathbf{x})$. As shown in Figure 2C, the learned score vectors align with normal vectors of the sphere (quantitative validation in C), replicating the theoretical prediction that score functions are normal to the data manifold (Stanczuk et al., 2024). Using these score functions, we compute geodesics and compare with linear interpolation. Figure 2D-F shows that while linear paths cut through the sphere (producing blurred intermediate images), our geodesics follow the manifold’s surface (maintaining high sample quality throughout). Similarly, for extrapolation, our approach follows the sphere’s curvature and preserves image structure, while linear extrapolation quickly departs from the manifold, causing severe distortions.
@@ 4.2 Rotated MNIST
To evaluate whether our score-based metric captures perceptually meaningful transformations, we conducted experiments on a modified version of the MNIST dataset where the transformation structure is well-defined. We created a custom Rotated MNIST dataset by systematically rotating each digit by angles ranging from $0^{\circ}$ to $350^{\circ}$ in $10^{\circ}$ increments, producing 36 rotated versions of each original digit. We then trained a denoising diffusion model on this dataset using a UNet-2D (Ronneberger et al., 2015) architecture with attention blocks. The model was trained for 100 epochs with a batch size of 256 (see more details in Appendix). Our hypothesis is that this trained model implicitly learns the data manifold, including the rotational structure of the dataset.
For interpolation experiments, we selected test set digit pairs with varying orientations, typically choosing the same digit rendered at different angles. Figure 3A compares our geodesic interpolation method against LERP, SLERP and Noise Diffusion. The results show that our geodesic approach produces trajectories that follow the rotational structure of the data manifold, creating smooth transformations that preserve digit identity while naturally rotating from one orientation to another. While LERP often produces unrealistic blending artifacts where intermediate frames superimpose features from both orientations rather than rotating continuously, our method follows the natural rotation transformation. Quantitatively, we measure the quality of interpolated frames using both PSNR and SSIM (Wang et al., 2004) metrics (Table 1), with our geodesic approach outperforming other methods in intermediate frame quality.
@@
PSNR $\uparrow$ 14.08 $\pm$ 0.12 13.64 $\pm$ 0.13 13.67 $\pm$ 0.08 14.98 $\pm$ 0.12
SSIM $\uparrow$ 0.578 $\pm$ 0.006 0.572 $\pm$ 0.007 0.568 $\pm$ 0.005 0.650 $\pm$ 0.006
@@ 4.3 Stable Diffusion & MorphBench
To demonstrate our method’s applicability to state-of-the-art diffusion models, we conducted experiments with Stable Diffusion 2.1 (Rombach et al., 2022), a powerful latent diffusion model trained on billions of image-text pairs. This allowed us to scale our approach on a more complex, higher-resolution image generation task than our previous experiments. Stable Diffusion operates in a compressed latent space rather than pixel space, making it an interesting test case for our score-based metric. The diffusion process occurs in a 4×64×64 latent space (representing 512×512 pixel, RGB images), where the data manifold has complex geometry reflecting natural image statistics. We apply our methodology by computing geodesics in this latent space using score estimates from the model’s denoising U-Net, then mapping the results back to image space using the model’s VAE decoder.
For evaluation, we utilize MorphBench (Zhang et al., 2024), a comprehensive benchmark for assessing image morphing capabilities. MorphBench consists of 90 diverse image pairs organized into two categories: (1) metamorphosis between different objects (66 pairs) and (2) animation of the same object with different attributes (24 pairs). This diversity allows us to evaluate our method across varying transformation complexities.
Figure 4 (see Appendix K, for more examples) presents qualitative comparisons between our geodesic method and baseline approaches (LERP, SLERP, and Noise Diffusion). Our method generates smoother, more natural-looking transitions that better preserve image coherence during the morphing process. While LERP and SLERP often produce blurry intermediate frames with unrealistic composites of both source and target images, and Noise Diffusion frequently generates images with artifacts, our geodesic interpolation creates a more natural progression by following meaningful paths on the data manifold.
Table 2 provides a comprehensive quantitative evaluation. Our geodesic method achieves the best performance on LPIPS (0.3582 vs. 0.3613 for LERP), a learned perceptual similarity metric that correlates strongly with human judgments of visual quality (Zhang et al., 2018). We also outperform all baselines on distribution-level metrics, with the lowest FID (140.61 vs. 148.19 for LERP) (Heusel et al., 2017) and KID scores (0.0863 vs. 0.0935 for LERP) (Bińkowski et al., 2018). These results indicate that our interpolated images are both perceptually coherent and maintain high sample quality that better matches the ”true” data distribution. While LERP achieves slightly better results on pixel-level metrics (PSNR: 21.28 vs. 20.88; SSIM: 0.6274 vs. 0.6180), it’s well-established that these metrics often fail to capture perceptual quality (Zhang et al., 2018), especially for high-resolution natural images.
@@
avg_SSIM $\uparrow$ 0.627 $\pm$ 0.012 0.575 $\pm$ 0.014 0.404 $\pm$ 0.012 0.618 $\pm$ 0.012
avg_PSNR $\uparrow$ 21.28 $\pm$ 0.18 20.36 $\pm$ 0.22 16.73 $\pm$ 0.20 20.88 $\pm$ 0.17
avg_LPIPS $\downarrow$ 0.361 $\pm$ 0.008 0.396 $\pm$ 0.008 0.048 $\pm$ 0.005 0.358 $\pm$ 0.007
FID $\downarrow$ 148.2 $\pm$ 5.0 171.5 $\pm$ 5.3 271.8 $\pm$ 4.7 140.6 $\pm$ 5.1
KID $\downarrow$ 0.094 $\pm$ 0.005 0.110 $\pm$ 0.006 0.186 $\pm$ 0.006 0.086 $\pm$ 0.005
@@ 5 Discussion
In this work, we introduced a score-based Riemannian metric derived from diffusion models that captures the intrinsic geometry of the data manifold without requiring explicit parameterization. Our approach leverages the Stein score function to define a metric that naturally adapts to the manifold structure, enabling geometric computations that respect the underlying data distribution. Through experiments on synthetic data, MNIST, and complex natural images, we demonstrated that geodesics computed using our score-based metric correspond to perceptually meaningful paths between images, outperforming conventional interpolation methods.
@@
An interesting pattern emerged when comparing performance across different datasets. On Rotated MNIST, our method outperformed baselines across pixel-level metrics (PSNR and SSIM). However, on the higher-resolution MorphBench dataset with Stable Diffusion, our method dominated in perceptual metrics (LPIPS) and distribution metrics (FID, KID) but was slightly outperformed by LERP on pixel-level metrics. This difference can be attributed to our geodesic computations for Stable Diffusion occurring in the compressed latent space rather than the pixel space. The VAE decoder introduces additional variability when mapping back to pixels, which may affect pixel-perfect reconstruction while still preserving—and even enhancing—perceptual quality. This observation aligns with the established understanding that pixel-level metrics often fail to capture perceptual similarity in high-resolution natural images, instead favoring blurry but pixel-wise accurate reconstructions over sharper, perceptually preferable images (Zhang et al., 2018).
@@
Limitations. Despite its strengths, our approach introduces additional computational complexity compared to direct methods like LERP or SLERP. Our implementation requires solving a non-convex optimization problem with several hundred gradient steps to converge, whereas alternative methods can be computed without any iterative optimization procedure. This reflects an open challenge in Riemannian optimization—efficiently computing geodesics on neural network-defined manifolds (Smith, 2014). As the field of Riemannian optimization advances, we anticipate future improvements that could significantly reduce this computational burden.
@@
A second limitation relates to the validation of extrapolation results. Although our experiments demonstrate that extrapolated images maintain high quality and follow natural extensions of geodesic paths, there are no established quantitative metrics to evaluate extrapolation quality in our settings. This makes objective comparisons challenging, though we view extrapolation primarily as a tool for exploring the implicit structure captured by diffusion models rather than a task requiring ground truth validation.
@@
Future Work. One exciting direction would be applying our approach to semantic image editing tasks (Brack et al., 2023; Zhang et al., 2023). By computing geodesic paths between edited and original images, our method could enable more natural and realistic transitions when performing attribute manipulations, style transfers, or object insertions.
@@
Additionally, the computational efficiency of geodesic calculation could be improved through techniques like neural surrogate models (Chen et al., 2018) that directly predict geodesic paths. Such models could enable real-time interactive editing tools that allow artists and designers to explore the learned manifold of a diffusion model while maintaining high sample quality throughout the editing process. Finally, our score-based metric provides a new lens for analyzing the geometry learned by diffusion models.
@@
Future work could explore using this geometric perspective to better understand model behavior, detect biases in learned distributions, or visualize how the data manifold evolves during training. By continuing to develop the connections between diffusion models and Riemannian geometry, we can gain deeper insights into how these powerful generative models represent complex data distributions.
Appendix A Related works
Riemannian Geometry in Diffusion Models. Recent advances have explored various approaches to construct metrics from generative models and navigating data manifolds. Stanczuk et al. (2024) demonstrated that diffusion models can encode the intrinsic dimension of data manifolds through score function analysis, establishing that score vectors point normal to the data manifold (Theorem 4.1). Building on these insights, Diepeveen et al. (2024) proposed a score-based pullback Riemannian metric that provides closed-form expressions for geodesics and distances on data manifolds. Similarly, Park et al. (2023) analyzed diffusion model latent spaces through pullback metrics derived from the encoding feature maps, enabling semantically meaningful image editing.
For manifold-aware navigation, several approaches have been developed. Kapusniak et al. (2024) proposed Metric Flow Matching (MFM), which learns interpolants by minimizing the kinetic energy of a data-induced Riemannian metric. Samuel et al. (2023) demonstrated that standard interpolation methods like LERP and SLERP produce suboptimal results in high-dimensional latent spaces, proposing norm-guided approaches with non-Euclidean metrics. Complementing interpolation methods, Yun et al. (2024) developed techniques for manifold extrapolation using diffusion model denoisers, showing that score functions can guide extrapolation while preserving manifold structure.
Related Metric Structures. The rank-1 perturbed metric $g(x)=I+\lambda s(x)s(x)^{T}$ represents a fundamental construction in differential geometry, with deep connections to deformation theory and general relativity. Llosa & Soler (2005) established that any 3-dimensional Riemannian metric can be locally expressed as $g=a\eta+\epsilon s\otimes s$ (Theorem 1), where $\eta$ is a constant curvature metric and $s$ is a differential 1-form. Coll (1999) generalize this result to arbitrary dimensions, establishing rank-1 perturbations as a canonical form for metric deformation in differential geometry. Remarkably, the Kerr-Schild class of metrics in general relativity meet this relation (Debney et al., 1969).
Beyond theoretical foundations, similar metric constructions have been applied in computational statistics and optimization contexts, and as in Hartmann et al. (2022) (see also (Hartmann et al., 2023; Bergamin et al., 2023; Yu et al., 2023a; b)), are termed the Monge metrics. Their construction embeds the target manifold into an extended space using scaled coordinates $\Xi(x)=(x,\alpha\log p(x))$, naturally inducing the metric $G(x)=I+\alpha^{2}\nabla\log p(x)\nabla\log p(x)^{T}$ for efficient MCMC sampling in parameter spaces. This approach creates position-dependent metrics that penalize movement along high-gradient directions, enabling more efficient exploration of complex probability distributions. Our construction follows the same fundamental mathematical principle, using $s(x)=\nabla\log p(x)$ to create directional anisotropy. However, rather than parameter space sampling, we apply this canonical metric form to data manifold exploration, where the score function provides normal directions that enable ambient space geodesic computation without explicit manifold parameterization.
Appendix B Proof of Riemannian Metric Properties
@@
To prove that $g(\bm{x})=\mathbf{I}+\lambda\cdot\bm{s}(\bm{x})\bm{s}(\bm{x})^{T}$ defines a valid Riemannian metric, we verify the fundamental properties:
Symmetry: Immediate since $(\bm{s}(\bm{x})\bm{s}(\bm{x})^{T})^{T}=\bm{s}(\bm{x})\bm{s}(\bm{x})^{T}$;
Positive-definiteness: For any non-zero vector $\bm{v}$:
$\bm{v}^{T}g(\bm{x})\bm{v}=\|\bm{v}\|^{2}+\lambda(\bm{s}(\bm{x})^{T}\bm{v})^{2}>0$
since $\lambda>0$ and $\|\bm{v}\|^{2}>0$ for $\bm{v}\neq\mathbf{0}$.
Thus $g(\bm{x})$ is a valid Riemannian metric on $\mathbb{R}^{N}$.
Appendix C Score vectors are Normal to the Data Manifold
Timestep Mean Angle (°) Angle Variance (°)
0 127.7 1.13
10 129.1 1.41
50 167.4 1.24
100 170.8 1.25
300 172.7 1.32
400 172.8 1.30
500 172.8 1.30
Table 3: Quantitative analysis of score-normal alignment across diffusion timesteps. Mean angle approaches 180° as timestep increases, indicating alignment with inward-pointing normals.
To validate the theoretical prediction (Stanczuk et al., 2024) that score vectors align with normal vectors of the data manifold, we implemented a comprehensive statistical analysis framework. For each timestep, we randomly sampled 10,000 points from our spherical dataset and computed the diffusion score function $\mathbf{s}(\mathbf{x})$ using our trained model. Since the data lies on a sphere, the ground truth normal vectors are the radial directions from the origin. We systematically analyzed the geometric relationship between scores and these normals by computing: (1) the angle between each score vector and the corresponding radial direction, (2) variance of the sampled score vectors.
As shown in Table 3, the mean angle between score and radial vectors rapidly approaches 173° as the timestep increases, demonstrating that scores strongly align with inward-pointing normals. The consistently small variance values indicate a tightly clustered distribution around the mean angle. This confirms that our approach, which uses $t=400$ in the main experiments, operates in a regime where score vectors reliably approximate normal vectors to the data manifold, with the consistent inward orientation expected from the maximum likelihood training of diffusion models. These results provide robust empirical validation for the geometric foundation of our metric deformation methodology.
Appendix D Parameters selection
D.1 Time $t$
In Appendix C we provide results for understanding when score vectors are normal to the data manifold. We build on this to select $t=400$, which is further supported by Bodin et al. (2025). The choice of $t\neq T$ has been reported multiple times in the literature (Mokady et al., 2023; Samuel et al., 2023).
D.2 Ablation study for $\lambda$
Since the true manifold structure for natural images is unknown, we developed a principled approach for selecting the penalty parameter $\lambda$ using controlled experiments on synthetic manifolds with known ground truth geodesics. On a 2-sphere with unitary radius embedded in $\mathbb{R}^{100}$, we studied how $\lambda$ affects geodesic approximation error by comparing our computed geodesics against analytical solutions (see Table 4) .
$\lambda$ Relative Error (%)
0 4.92
10 0.50
100 0.11
1000 0.07
5000 0.07
10000 0.07
Table 4: Performance comparison across different $\lambda$ values
The error decays approximately as $A\cdot\exp(-\lambda/\lambda_{0})+B$ (5). Similar results on a 50-sphere in $\mathbb{R}^{100}$ confirmed $\lambda=1000$ achieves $\sim$0.07% error consistently and after that there are no more improvements. We used this same value for all experiments in the paper.
Figure 5: Relative error as a function of penalty parameter $\lambda$. The dashed red line shows the exponential fit $A\exp(-\lambda/\lambda_{0})+B$.
D.3 Extrapolation Hyperparameters $\beta$ and $\varepsilon$
For $\beta$ we sticked to the default value (i.e. 0.9) used in Adam (Kingma, 2014) and Riemannian Adam (Becigneul & Ganea, 2019), as the idea of using a moving-average is similar in principle. While for $\varepsilon$, we stick to 0.1, giving more importance to the momentum vector. We consider the extrapolation results more as a proof-of-principle and a qualitative results that could enable future works.
Appendix E Geodesic Algorithm Details
E.1 Discretization and Numerical Integration